OpenStudy (anonymous):

College Algebra: Can anyone help me?? If $2500 is invested in the account that pays interest compounded continuously at a rate of 4%, how long will it take to double the investment? --Formula: A=Pe^rt --Answer should be 17.3 years --HOW do I get the answer 17.3? Can someone show me?

OpenStudy (danjs):

P=2500 r=0.04 A = 2*2500 = 5000

OpenStudy (danjs):

solve for the only thing left, time t

OpenStudy (anonymous):

Oh my dear lord, thank you so much!!! I been trying to solve this problem for over 30 minutes now. Thank you thank you thank you!!! :)

OpenStudy (danjs):

you get it?

OpenStudy (anonymous):

Yessss I do!! :)

OpenStudy (danjs):

2 = e^(0.04t)

OpenStudy (danjs):

cool

OpenStudy (anonymous):

Yes I solved it

OpenStudy (anonymous):

Thank you so so so very much!! :)